Suppose that the world price for a good is 40, and the domestic demand and supply curves are given by the following equations:

Demand: P=80-2Q

Supply: P=5+3Q

a. (2 points) How much is consumed? How much is produced domestically?

b. (2 points) What are the values of producer and consumer surplus?

c. (2 points) If a 10% tariff is imposed, by how much do consumption and domestic

production change?

d. (2 points) What is the change in consumer and producer surplus?

e. (2 points) How much revenue does the government earn from the tariff?


1
Expert's answer
2021-04-26T19:34:24-0400

a. The amount consumed is:

80 - 2Q = 40,

Q = 20 units.

The amount produced domestically is:

5 + 3Q = 40,

Q = 11.67 units.

b. The values of producer and consumer surplus are:

PS = 0.5×11.67×(40 - 5) = 204.23,

CS = 0.5×20×(80 - 40) = 400.

c. If a 10% tariff is imposed, then consumption will decrease to 18 units and domestic production will increase to 13 units.

d. Consumer surplus will decrease and producer surplus will increase.

e. The government will earn: 40×0.1×(18 - 13) = 20.